5-13. Bill Mitselfik has purchased a bond that was issued by Acme Chemical. This bond has a face value of $1,000 and pays a dividend of 8% per year, compounded semi-annually. Bill bought the bond three years ago at face value and there are seven years remaining until the bond matures. Bill wishes to sell it now for a price that will result in Bill earning an annual yield of 10% compounded semi-annually. What price does Bill need to sell the bond for to earn his desired return?please use tables not excel
Answer #1

Semi-annual interest on bond = $1,000 x 8% x (1/2) = $40

Semi-annual yield = 10% x (1/2) = 5%

Number of compounding periods = 7 x 2 = 14

Bond price ($) = Present value of future interest payments + Present value of redemption price (face value)

= 40 x P/A(5%, 14) + 1,000 x P/F(5%, 14)

= 40 x 9.8986** + 1,000 x 0.5051**

= 395.94 + 505.1

= 901.04

**From P/A and P/F Factor tables